# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

### v1

@@ 2.1.2. Right Hilbert $C(X)$-modules
When $A=C(X)$, we can construct Hilbert $A$-modules from vector bundles. For a good introduction to the theory of bundles, see [20]. Let $X$ be a compact metric space and $\mathscr{V}=[T,p,X]$ a vector bundle of rank $n$. (Here $T$ denotes the total space of the bundle, $X$ the base space, and $p:T\to X$ the projection map.) We denote the set of continuous sections of $\mathscr{V}$ by $\Gamma(\mathscr{V})$. Then $\Gamma(\mathscr{V})$ has a natural right $C(X)$-module structure given by multiplication of a section by a continuous function. n fact, it is a right Hilbert $C(X)$-module: Let $\mathcal{A}=\{(U_{j},h_{j})\}_{j=1}^{N}$ be an atlas of $\mathscr{V}$ and let $\gamma_{1},\dots,\gamma_{N}$ a partition of unity subordinate to $U_{1},\dots,U_{N}$. Then the following $C(X)$-valued inner product
$\langle\xi,\eta\rangle_{\Gamma(\mathscr{V})}(x):=\sum_{j=1}^{N}\gamma_{j}(x)\langle h_{j}^{-1}(\xi(x)),h_{j}^{-1}(\eta(x))\rangle_{\mathbb{C}^{n}}$ (1)
makes $\Gamma(\mathscr{V})$ into a right Hilbert $C(X)$-module. Note that the inner product given in (1) does not depend on the choice of atlas. Moreover, with pairwise orthonormal vectors $s_{1},\dots,s_{n}$ of $\mathbb{C}^{n}$, the continuous sections given by
@@
$\xi=\sum_{j=1}^{N}\sum_{k=1}^{n}\xi_{j,k}\langle\xi_{j,k},\xi\rangle_{\Gamma(\mathscr{V})}$
@@
By the Serre–Swan theorem [37, Theorem 2] every finitely-generated projective right $C(X)$-module is of the form $\Gamma(\mathscr{V})$ for a vector bundle over $X$. Furthermore, if $\mathcal{E}$ is a right Hilbert $C(X)$-module, there is a unitary isomorphism $U:\mathcal{E}\to\Gamma(\mathscr{V})$, where $\Gamma(\mathscr{V})$ is equipped with an inner product as defined in (1) with respect to any choice of atlas for $\mathscr{V}$.
@@ 2.1.3. Homeomorphisms twisted by line bundles
We recall from [2] that, when a Hilbert $C(X)$-module is left and right full, it always arises from a homeomorphism twisted by a line bundle over $X$.
Let $\alpha:X\to X$ be a homeomorphism and let $\mathscr{V}=[T,p,X]$ be a line bundle. We denote the pullback bundle by $\alpha^{*}\mathscr{V}$: the fibre $(\alpha^{*}\mathscr{V})_{x}$ of $\alpha^{*}\mathscr{V}$ at $x\in X$ is the fibre $\mathscr{V}_{\alpha(x)}$ of $\mathscr{V}$ at $\alpha(x)$, and $(x,\lambda)\mapsto h_{\alpha(U)}(\alpha(x),\lambda)$, $\lambda\in\mathbb{C}$, is a chart of $\alpha^{*}\mathscr{V}|_{U}$ whenever $(\alpha(U),h_{\alpha(U)})$ is a chart of $\mathscr{V}|_{\alpha(U)}$.
@@
${}_{C(X)}\langle\xi_{1},\xi_{2}\rangle=\langle\xi_{2},\xi_{1}\rangle_{C(X)}\circ\alpha^{-1}.$
@@
Suppose that $\mathcal{E}$ is a non-zero Hilbert $C(X)$-bimodule which is finitely-generated and projective as a right Hilbert $C(X)$-module. Then there exist a compact subset $Y\subset X$, line bundle $\mathscr{V}$ over $X$ and homeomorphisms $\alpha:X\to Y$, $\beta:Y\to X$ such that
$\mathcal{E}_{C(X)}\cong\Gamma(\mathscr{V}),\quad_{C(X)}\mathcal{E}\cong\Gamma(\beta^{*}\mathscr{V}),$
and $f\xi=\xi f\circ\alpha$ for every $f\in C(X)$ and every $\xi\in\mathcal{E}$. Moreover, if $\mathcal{E}$ is left full, then we may take $X=Y$ and $\mathcal{E}\cong\Gamma(\mathscr{V},\alpha)$.
@@ 2.1.4. Orbit-breaking bimodules $\mathcal{E}_{Y}$
The main Hilbert bimodules of focus of this paper are the orbit-breaking subbimodules $\mathcal{E}_{Y}$ of $\mathcal{E}=\Gamma(\mathscr{V},\alpha)$, which we define below.
@@ 2.2. Cuntz–Pimsner algebras
Cuntz–Pimsner algebras were originally introduced by Pimsner [33] and subsequently generalized by Katsura [22]. The class of Cuntz–Pimsner algebras includes all Cuntz–Krieger algebras, crossed products by $\mathbb{Z}$, crossed products by partial homeomorphisms and several other naturally occurring $\mathrm{C}^{*}$-algebras, see [22, Section 3].
Definition 2.4 ([23, Definition 2.1]).
@@
We say that the representation $(\pi,\tau)$ is covariant [23, Definition 3.4] if in addition
@@
    $\pi(a)=\psi_{\tau}(\varphi_{\mathcal{E}}(a))$, for every $a\in J_{\mathcal{E}}:=\varphi_{\mathcal{E}}^{-1}(\mathcal{K}(\mathcal{E}))$.
In [23, Section 4] Kastura constructs a universal covariant representation $(\pi_{u},\tau_{u})$ satisfying the following: for any covariant representation $(\pi,\tau)$ of $\mathcal{E}$, there exists a surjective -homomorphism $\rho\colon\mathrm{C}^{*}(\pi_{u},\tau_{u})\rightarrow\mathrm{C}^{*}(\pi,\tau)$ such that $\pi=\rho\circ\pi_{u}$ and $\tau=\rho\circ\tau_{u}$, where ${\mathrm{C}}^{*}(\pi,\tau)$ is the $\mathrm{C}^{*}$-subalgebra of $B$ generated by $\pi(A)$ and $\tau(\mathcal{E})$.
Definition 2.5 ([23, Definition 3.5], see also [1, Definition 2.4]).
@@
A representation $(\pi,\tau)$ of $\mathcal{E}$ is said to admit a gauge action if, for each $z\in\mathbb{T}$, there exists a -homomorphism $\beta_{z}:\mathrm{C}^{*}(\pi,\tau)\to\mathrm{C}^{*}(\pi,\tau)$ such that $\beta_{z}(\pi(a))=\pi(a)$ and $\beta_{z}(\tau(\xi))=z\tau(\xi)$ for every $a\in A$ and every $\xi\in\mathcal{E}$ [23, Definition 5.6]. The gauge-invariant uniqueness theorem for Cuntz–Pimsner algebras is stated as follows.
Theorem 2.7 ([23, Theorem 6.4]).
@@
$\mathcal{O}(\mathcal{E})=\overline{\bigoplus_{n\in\mathbb{Z}}}$
@@
where $\hat{\xi}:\mathcal{E}\to A$ is an $A$-bilinear map defined by $\hat{\xi}(\eta)=\langle\xi,\eta\rangle_{\mathcal{E}}$, $\xi,\eta\in\mathcal{E}.$ Then $\hat{\xi}\in\mathcal{E}^{\dagger}$, the dual Hilbert $A$-bimodule consisting of bounded $A$-linear maps $\varphi:\mathcal{E}\to A$, with actions given by
@@
When $\mathcal{E}$ is finitely-generated projective as both a left and right $A$-module, we have that =^:=$\mathcal{E}^{\dagger}=\widehat{\mathcal{E}}:=\{\hat{\xi}\mid\xi\in\mathcal{E}\}$, see the discussion following Theorem 1.3 in [30]. We make $\widehat{\mathcal{E}}$ into a Hilbert $A$-bimodule by defining
@@
Note that if $\mathcal{E}^{\dagger}\cong\widehat{\mathcal{E}}$, then $\mathcal{E}$ is invertible in the sense that
@@ 3. Tensor products of Hilbert $C(X)$-bimodules $\Gamma(\mathscr{V},\alpha)$
As tensor products of Hilbert $C(X)$-bimodules play an important role in the construction of the relevant Cuntz–Pimsner algebras, we discuss the relation between the tensor products of Hilbert $C(X)$-bimodules and the Hilbert $C(X)$-bimodule constructed from the tensor product of the underlying line bundles and the composition of homeomorphisms.
@@
Whenever $\mathscr{V}|_{\alpha^{i}(U)}$ is trivial for all $i=0,1,\dots,n-1$ (via $h_{\alpha^{i}(U)}$), we have a chart map
$(x,\lambda)\mapsto\lambda h_{\alpha^{n-1}(U)}(\alpha^{n-1}(x),1)\otimes\cdots\otimes h_{U}(x,1)$
of the vector bundle $\mathscr{V}^{(n)}$ over $U$.
@@
$=$ $\,\langle\xi_{2}\otimes\cdots\otimes\xi_{n},\eta_{2}\otimes\cdots\otimes\eta_{n}\rangle_{\mathcal{E}^{\otimes n-1}}\langle\xi_{1},\eta_{1}\rangle_{\mathcal{E}}\circ\alpha^{n-1}(x)$
@@
Let $\eta_{j}$ be the generators of $\mathcal{E}=\Gamma(\mathscr{V})$ as in (2). From the definition of $\mathscr{V}^{(n)}$, one can show that the sections $\psi(\eta_{i_{1}}\otimes\eta_{i_{2}}\otimes\cdots\otimes\eta_{i_{n}})$ generate $\Gamma(\mathscr{V}^{(n)},\alpha^{n})$. Then for $\xi\in\Gamma(\mathscr{V},\alpha)^{\otimes n}$, the element
@@
Let $\alpha:X\to X$ be a homeomorphism and $\mathscr{V}$ be a line bundle on $X$. Let $\psi:\mathcal{E}^{\otimes m}\to\Gamma(\mathscr{V}^{(m)},\alpha^{m})$ be the isomorphism given by Proposition 3.1. Let $Y\subset X$ be a non-empty closed subset and denote by $\mathcal{E}_{Y}$ the associated orbit-breaking bimodule. Then for any element $\xi\in\mathcal{E}^{\otimes m}$, we have that $\xi\in\mathcal{E}_{Y}^{\otimes m}:=(\mathcal{E}_{Y})^{\otimes m}$ if and only if $\psi(\xi)$ vanishes on $\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. If $m=0$, then this set is empty and by $\psi$ we mean the identity map of $C(X)$.
@@
$=\eta_{1}\otimes\cdots\otimes\eta_{m}(f_{1}\circ\alpha^{m})\cdots(f_{m-1}\circ\alpha^{2})(f_{m}\circ\alpha),$
thus $\psi(\xi)(x)=0$ whenever $x\in\alpha^{-1}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Since such $\xi$ generate $\mathcal{E}_{Y}^{\otimes m}$, the only if part follows.
Conversely, suppose $\xi\in\mathcal{E}^{\otimes m}$ satisfies $\psi(\xi)(x)=0$ for every $x\in\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Let $\{h_{j}:U_{j}\times\mathbb{C}\to\mathscr{V}|_{U_{j}}\}_{j=1,\ldots,n}$ be an atlas for $\mathscr{V}$ and $\gamma_{1},\ldots,\gamma_{n}$ a partition of unity subordinate to $U_{1},\ldots,U_{n}$. Then $\eta_{j}$, $1\leq j\leq n$, given by $\eta_{j}(x)=h_{j}(x,{\gamma_{j}}^{1/2}(x))$, generates $\mathcal{E}$ as a right Hilbert $C(X)$-module, hence $\eta_{i_{1}}\otimes\cdots\otimes\eta_{i_{m}}$, $1\leq i_{1},\dots,i_{m}\leq n$ generates $\mathcal{E}^{\otimes m}$. Then the functions
@@
vanishes on $\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Since the functions $f_{i_{1},i_{2},\ldots,i_{m}}$ are linear combinations of positive ones, assuming that they are positive in the following computation, we have

### v2

@@ 2.1.2. Right Hilbert $C(X)$-modules
When $A=C(X)$, we can construct Hilbert $A$-modules from vector bundles. For a good introduction to the theory of bundles, see [22]. Let $X$ be a compact metric space and $\mathscr{V}=[T,p,X]$ a Hermitian vector bundle of rank $n$. (Here $T$ denotes the total space of the bundle, $X$ the base space, and $p:T\to X$ the projection map.) We denote the set of continuous sections of $\mathscr{V}$ by $\Gamma(\mathscr{V})$. Then $\Gamma(\mathscr{V})$ has a natural right $C(X)$-module structure given by multiplication of a section by a continuous function. In fact, it is a right Hilbert $C(X)$-module: Let $\mathcal{A}=\{(U_{j},h_{j})\}_{j=1}^{N}$ be an atlas of $\mathscr{V}$ and let $\gamma_{1},\dots,\gamma_{N}$ a partition of unity subordinate to $U_{1},\dots,U_{N}$. Chart maps are always assumed to preserve inner products: $\langle h_{j}(x,v),h_{j}(x,w)\rangle_{p^{-1}(x)}=\langle v,w\rangle_{\mathbb{C}^{n}}$. Then the $C(X)$-valued inner product
$\langle\xi,\eta\rangle_{\Gamma(\mathscr{V})}(x):=\sum_{j=1}^{N}\gamma_{j}(x)\langle h_{j}^{-1}(\xi(x)),h_{j}^{-1}(\eta(x))\rangle_{\mathbb{C}^{n}},\quad\xi,\eta\in\Gamma(\mathscr{V}),$ (1)
makes $\Gamma(\mathscr{V})$ into a right Hilbert $C(X)$-module. Note that the inner product given in (1) does not depend on the choice of atlas nor the choice of a partition of unity. Moreover, with pairwise orthonormal vectors $s_{1},\dots,s_{n}$ of $\mathbb{C}^{n}$, the continuous sections given by
@@
$\xi=\sum_{j=1}^{N}\sum_{k=1}^{n}\xi_{j,k}\langle\xi_{j,k},\xi\rangle_{\Gamma(\mathscr{V})},$
@@
By the Serre–Swan theorem [40, Theorem 2] every finitely generated projective right $C(X)$-module is of the form $\Gamma(\mathscr{V})$ for a vector bundle over $X$. Furthermore, if $\mathcal{E}$ is a right Hilbert $C(X)$-module, there is a unitary isomorphism $U:\mathcal{E}\to\Gamma(\mathscr{V})$, where $\Gamma(\mathscr{V})$ is equipped with an inner product as defined in (1) with respect to any choice of atlas for $\mathscr{V}$.
@@ 2.1.3. Homeomorphisms twisted by line bundles
We recall from [2] that, when a Hilbert $C(X)$-bimodule is left and right full, it always arises from a homeomorphism twisted by a line bundle over $X$.
Let $\alpha:X\to X$ be a homeomorphism and let $\mathscr{V}=[T,p,X]$ be a line bundle. We denote the pullback bundle by $\alpha^{*}\mathscr{V}$: the fibre $(\alpha^{*}\mathscr{V})_{x}$ of $\alpha^{*}\mathscr{V}$ at $x\in X$ is the fibre $\mathscr{V}_{\alpha(x)}$ of $\mathscr{V}$ at $\alpha(x)$.
@@
${}_{\Gamma(\mathscr{V},\alpha)}\langle\xi_{1},\xi_{2}\rangle=\langle\xi_{2},\xi_{1}\rangle_{\Gamma(\mathscr{V})}\circ\alpha^{-1}.$
@@
Let $\mathcal{E}$ be a non-zero Hilbert $C(X)$-bimodule. Suppose that $\mathcal{E}$ is finitely generated and projective as a right Hilbert $C(X)$-module. Then there exist a compact metric space $Y\subset X$, line bundle $\mathscr{V}=[T,p,X]$ and homeomorphisms $\alpha:X\to Y$, $\beta:Y\to X$ such that $\mathcal{E}_{C(X)}\cong\Gamma(\mathscr{V})$, ${}_{C(X)}\mathcal{E}\cong\Gamma(\beta^{*}\mathscr{V})$ and $f\xi=\xi f\circ\alpha$ for every $f\in C(X)$ and every $\xi\in\mathcal{E}$. If $\mathcal{E}$ is left full, then we may take $X=Y$ and $\mathcal{E}\cong\Gamma(\mathscr{V},\alpha)$.
@@ 2.1.4. Orbit-breaking bimodules $\mathcal{E}_{Y}$
The main objects of focus of this paper are the orbit-breaking subbimodules $\mathcal{E}_{Y}$ of $\mathcal{E}=\Gamma(\mathscr{V},\alpha)$, which we define below.
@@ 2.2. Cuntz–Pimsner algebras
Cuntz–Pimsner algebras were originally introduced by Pimsner [36] and subsequently generalized by Katsura [24]. The class of Cuntz–Pimsner algebras includes all Cuntz–Krieger algebras, crossed products by $\mathbb{Z}$, crossed products by partial homeomorphisms and several other naturally occurring $\mathrm{C}^{*}$-algebras, see [24, Section 3].
Definition 2.4 ([25, Definition 2.1]).
@@
We say that the representation $(\pi,\tau)$ is covariant [25, Definition 3.4] if in addition
@@
    $\pi(a)=\psi_{\tau}(\varphi_{\mathcal{E}}(a))$, for every $a\in J_{\mathcal{E}}:=\varphi_{\mathcal{E}}^{-1}(\mathcal{K}(\mathcal{E}))\cap(\ker\varphi_{\mathcal{E}})^{\perp}.$
In [25, Section 4] Katsura constructs a universal covariant representation $(\pi_{u},\tau_{u})$ satisfying the following: for any covariant representation $(\pi,\tau)$ of $\mathcal{E}$, there exists a surjective -homomorphism $\rho\colon\mathrm{C}^{*}(\pi_{u},\tau_{u})\rightarrow\mathrm{C}^{*}(\pi,\tau)$ such that $\pi=\rho\circ\pi_{u}$ and $\tau=\rho\circ\tau_{u}$, where ${\mathrm{C}}^{*}(\pi,\tau)$ is the $\mathrm{C}^{*}$-subalgebra of $B$ generated by $\pi(A)$ and $\tau(\mathcal{E})$.
Definition 2.5 ([25, Definition 3.5], see also [1, Definition 2.4]).
@@
A representation $(\pi,\tau)$ of $\mathcal{E}$ is said to admit a gauge action if, for each $z\in\mathbb{T}$, there exists a -homomorphism $\beta_{z}:\mathrm{C}^{*}(\pi,\tau)\to\mathrm{C}^{*}(\pi,\tau)$ such that $\beta_{z}(\pi(a))=\pi(a)$ and $\beta_{z}(\tau(\xi))=z\tau(\xi)$ for every $a\in A$ and every $\xi\in\mathcal{E}$ [25, Definition 5.6]. The gauge-invariant uniqueness theorem for Cuntz–Pimsner algebras is stated as follows.
Theorem 2.7 ([25, Theorem 6.4]).
@@
$\mathcal{O}(\mathcal{E})=\overline{\bigoplus_{n\in\mathbb{Z}}E_{n}}$
@@
where $\hat{\xi}:\mathcal{E}\to A$ is an $A$-bilinear map defined by $\hat{\xi}(\eta)=\langle\xi,\eta\rangle_{\mathcal{E}}$, $\xi,\eta\in\mathcal{E}.$ Then $\hat{\xi}\in\mathcal{E}^{\dagger}$, the dual Hilbert $A$-bimodule consisting of bounded $A$-linear maps $\phi:\mathcal{E}\to A$, with actions given by
@@
When $\mathcal{E}$ is finitely generated projective as both a left and right $A$-module, we have that $\mathcal{E}^{\dagger}=\widehat{\mathcal{E}}$, see the discussion following Theorem 1.3 in [33]. We make $\widehat{\mathcal{E}}$ into a Hilbert $A$-bimodule by defining
@@
Note that if $\mathcal{E}^{\dagger}=\widehat{\mathcal{E}}$, then $\mathcal{E}$ is invertible in the sense that
@@ 3. Tensor products of Hilbert $C(X)$-bimodules $\Gamma(\mathscr{V},\alpha)$
As we will use the tensor products of line bundles in the construction of the building blocks of the RSH decomposition of an orbit-breaking algebra, we discuss the relation between the tensor products of Hilbert $C(X)$-bimodules and the Hilbert $C(X)$-bimodule constructed from the tensor product of the underlying line bundles and the composition of homeomorphisms.
@@
Whenever $\mathscr{V}|_{U_{i}}$, $\alpha^{i}(x)\in U_{i}$, is trivial via $h_{U_{i}}$ for all $i=0,\dots,n-1$, we have a chart map
$(x,\lambda)\mapsto\lambda\,h_{U_{n-1}}(\alpha^{n-1}(x),1)\otimes\cdots\otimes h_{U_{1}}(\alpha(x),1)\otimes h_{U_{0}}(x,1)$
of the vector bundle $\mathscr{V}^{(n)}$ over $\cap_{i=0}^{n-1}\alpha^{-i}(U_{i})$.
@@
$=\langle\xi_{2}\otimes\cdots\otimes\xi_{n},\eta_{2}\otimes\cdots\otimes\eta_{n}\rangle_{\mathcal{E}^{\otimes n-1}}(x)\langle\xi_{1},\eta_{1}\rangle_{\mathcal{E}}\circ\alpha^{n-1}(x)$
@@
Let $\eta_{j}$ be the generators of $\mathcal{E}=\Gamma(\mathscr{V})$ as in (2). From the definition of $\mathscr{V}^{(n)}$, one can show that the sections $\psi(\eta_{i_{1}}\otimes\eta_{i_{2}}\otimes\cdots\otimes\eta_{i_{n}})$ generate $\Gamma(\mathscr{V}^{(n)},\alpha^{n})$. Then for $\xi\in\Gamma(\mathscr{V}^{(n)},\alpha^{n})$, the element
@@
Let $\alpha:X\to X$ be a homeomorphism and $\mathscr{V}$ be a line bundle on $X$. Let $\psi:\mathcal{E}^{\otimes m}\to\Gamma(\mathscr{V}^{(m)},\alpha^{m})$ be the isomorphism given by Proposition 3.1 with $\mathcal{E}=\Gamma(\mathscr{V},\alpha)$. Let $Y\subset X$ be a non-empty closed subset and denote by $\mathcal{E}_{Y}$ the associated orbit-breaking bimodule. Then for any element $\xi\in\mathcal{E}^{\otimes m}$, we have that $\xi\in\mathcal{E}_{Y}^{\otimes m}:=(\mathcal{E}_{Y})^{\otimes m}$ if and only if $\psi(\xi)$ vanishes on $\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. If $m=0$, then this set is empty and by $\psi$ we mean the identity map of $C(X)$.
@@
$=\eta_{1}\otimes\cdots\otimes\eta_{m}(f_{1}\circ\alpha^{m})\cdots(f_{m-1}\circ\alpha^{2})(f_{m}\circ\alpha).$
Thus $\psi(\xi)(x)=0$ whenever $x\in\alpha^{-1}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Since such $\xi$ generate $\mathcal{E}_{Y}^{\otimes m}$, the “only if” part follows.
Conversely, suppose $\xi\in\mathcal{E}^{\otimes m}$ satisfies $\psi(\xi)(x)=0$ for every $x\in\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Let $\{h_{j}:U_{j}\times\mathbb{C}\to\mathscr{V}|_{U_{j}}\}_{j=1,\ldots,n}$ be an atlas for $\mathscr{V}$ and $\gamma_{1},\ldots,\gamma_{n}$ a partition of unity subordinate to $U_{1},\ldots,U_{n}$. Then $\eta_{j}$, $1\leq j\leq n$, given by $\eta_{j}(x)=h_{j}(x,{\gamma_{j}}^{1/2}(x))$, generate $\mathcal{E}$ as a right Hilbert $C(X)$-module, and so $\eta_{i_{1}}\otimes\cdots\otimes\eta_{i_{m}}$, $1\leq i_{1},\dots,i_{m}\leq n$, generates $\mathcal{E}^{\otimes m}$. The functions
@@
vanish on $\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Since the functions $f_{i_{1},i_{2},\ldots,i_{m}}$ are linear combinations of positive ones, we may assume that they are positive. Then we have